OpenStudy (anonymous):

A mountain has a rise of 8,480 feet from its base. If there is a 2 degree drop in temperature for every rise of 530 feet in altitude, how many degrees difference is there between the ground temperature and the temperature at the peak? (Points : 1) 32 1060 16 4240 I'm guessing 32...but not sure.

OpenStudy (anonymous):

\[\Delta T= \frac{8480}{530} \times 2 = 32\]

OpenStudy (anonymous):

so is that the correct answer? thats the same way i did it

OpenStudy (anonymous):

At top altitude = 8480 530 feet in altitude,temperature falls by 2 degrees. Therefore, 8480/530 = 16 16*2 = 32 The difference between ground and peak of mountain temperatures is 32 degrees.